OpenStudy (anonymous):

Write the equation of the line which passes through (1, 2) and with slope of 13 in standard form.

OpenStudy (anonymous):

Easy way to get a medal!!!!

OpenStudy (anonymous):

slope of -1/3 not 13

OpenStudy (texaschic101):

y = mx + b slope(m) = -1/3 (1,2)..x = 1 and y = 2 now we sub 2 = -1/3(1) + b 2 = -1/3 + b 2 + 1/3 = b 6/3 + 1/3 = b 7/3 = b y = -1/3x + 7/3 -- multiply entire equation by 3 3y = -x + 7 -- add x to both sides x + 3y = 7 <== standard form

OpenStudy (anonymous):

thx